OpenStudy (anonymous):

Complex numbers question! One of the roots of z is a=cos2-isin2. Find the other roots and trace them on the graph.

OpenStudy (anonymous):

idk, sowy

OpenStudy (anonymous):

z has fourth degree, so the solutions are four

OpenStudy (nipunmalhotra93):

clearly z lies on the unit circle. Also, as we are finding fourth roots, the arguments of two successive roots will differ by Pi/2 on the circle. You are already given that one root is e^(-2i), so the other three roots are, \[e ^{i (-2+\pi/2)},e ^{i (-2+\pi)},e ^{i (-2+3\pi/2)}\]

OpenStudy (anonymous):

Can you explain to me why the arguments differ by pi/2?

OpenStudy (anonymous):

And why is it e^(-2i) and not e^(2i)?

OpenStudy (nipunmalhotra93):

you must be aware of the fact that for any complex number on the unit circle (in the complex plane of course), the nth roots form a regular n-gon inscribed in the unit circle. This also means that the angle subtended at the centre of the circle by each side of the n-gon will be equal. So, as the there are n sides, the measure of each angle will be 2Pi/n. Here, n=4. So, the roots form a square. So, angle subtended at the centre by each side of the square will be 2Pi/4 which is Pi/2. So, the difference of arguments between two successive roots will be Pi/2. You have cos2-i sin2. e^(ix)=cosx+i sinx. So, e^(-ix)=cosx+i sin(-x)=cosx-i sin(x)

OpenStudy (anonymous):

Ok i get it. In addition is it okay if I reason by adding as a solution the conjugate number of the given root?

OpenStudy (anonymous):

Thanks you by the way.

OpenStudy (anonymous):

On my book the results are: sin2-icos2 -cos2+isin2 sin2+icos2

OpenStudy (nipunmalhotra93):

@naylah no you can't add a conjugate. That's because you're finding the root of a non real number. It'd have been fine had it been a real number. and you're welcome :)

OpenStudy (anonymous):

But why aren't the results similar to the ones of my book?

OpenStudy (anonymous):

My professor is very strict about the method so he really wants stuff done his way...

OpenStudy (nipunmalhotra93):

your first answer is not correct. are you sure that's not a typo? it should be -sin2-icos2 instead of sin2-icos2

OpenStudy (anonymous):

yeah I forgot a sign, but how do I manage to come this far?